OpenStudy (curry):

DETERMINE 2 PAIRS OF POLAR COORDINATES FOR THE POINT (((-3RAD3)/2),(3/2)) WITH O<=THETA<=2PI

Directrix (directrix):

Is this point given in Cartesian and you want to convert it to polar?

OpenStudy (curry):

no just a normal coordinate plane

Directrix (directrix):

That's what Cartesian is, the x-y plane. I need to know if x = (-3RAD3)/2) and y = 3/2

OpenStudy (curry):

ye

OpenStudy (curry):

i figured out an answer but it is conflicting with other people's

OpenStudy (curry):

???? u there?

Directrix (directrix):

I was working with another person when you pinged me or whatever. I'm here now.

Directrix (directrix):

Who are these other people? Maybe they are correct.

OpenStudy (curry):

idk

OpenStudy (curry):

they gave me the answers (3, -60) (3,300) which is what i got but the right answer in the book is (3, 120) (2,-240)

Directrix (directrix):

You did not specify degrees so I chose radians.

OpenStudy (curry):

thats fine but can u elaborate on what u jusst said

Directrix (directrix):

The point is in the second quadrant with theta = 120 and r = 3/2. ( 3/2, 120) in polar form. How about ( -3/2, -60) and (3/2, 120) for one pair. For another, (3/2, 480) and (-3/2, -60)

Directrix (directrix):

Did you check these points? I'd like to know what the others got because there are infinitely many possibilities.

OpenStudy (curry):

well they gave me -60 and 300 for the angles but i still dont get how u can get 120

Directrix (directrix):

(3, 120) (2,-240) ==> the r values cannot be different.

OpenStudy (curry):

but aren't they different 3 and 2

OpenStudy (curry):

or is that typo?

Directrix (directrix):

Look at this restriction we ignored --> WITH O<=THETA<=2PI

Directrix (directrix):

I think you typo-ed.